A two sector economy has a total income of rs 150 billion and marginal propensity to consume (MPC) is 66.67%. How much does this economy needed to invest once for all to increase it total income by 100%.

1
Expert's answer
2021-11-21T16:52:18-0500

As we have Y= 150billion 

MPC = 66.67% or 0.67 therefore MPS will be =1-MPC and MPS= 0.33 

Because MPS+MPC=1

Investment multiplier (k) = 1/MPS or ∆Y/∆I.

k = 1/0.33 which is 3.03 

 We know that k = 3.03 we can use this to find investment needed to increase 100%income level .

100%invome level is itself it means ∆Y=150billion and k=3.03 

 

k= ∆Y/∆I : 3.03= 150/∆I 

∆I= 49.5billion 


To increase the 100%income level with MPC= 0.67 investment must be 49.5billion .